Let A= {1,2,3,4,5,} and R be a relation on A, such that
R = [(1,4), (2,2), (3,4), (3,5), (4,1), (5,2), (5,5)]. Use warshall’s Algorithm to find the matrix of transitive closure of R
1
Expert's answer
2020-11-02T20:34:39-0500

"R = \{(1,4), (2,2), (3,4), (3,5), (4,1), (5,2), (5,5)\}"


Here are the steps of the Warshall’s algorithm:


Step 1. Assign initial values "W=M_R, k=0".


Step 2. Execute "k:=k+1."


Step 3. For all "i\ne k" such that "w_{ik}=1", and for all "j" execute the operation "w_{ij}=w_{ij}\lor w_{kj}."


Step 4. If "k=n", then stop: we have the solution "W=M_{R^*}", else go to the step 2.



"n=|A|=5."


"W^{(0)}=M_R=\left(\begin{array} {cccccc} 0 & 0 & 0 & 1 & 0 \\ 0 & 1 & 0 & 0 & 0 \\ 0 & 0 & 0 & 1 & 1 \\ 1 & 0 & 0 & 0 & 0 \\ 0 & 1 & 0 & 0 & 1 \end{array}\right)"



"W^{(1)}=\left(\begin{array} {cccccc} 0 & 0 & 0 & 1 & 0 \\ 0 & 1 & 0 & 0 & 0 \\ 0 & 0 & 0 & 1 & 1 \\ 1 & 0 & 0 & 1 & 0 \\ 0 & 1 & 0 & 0 & 1 \end{array}\right)"



"W^{(2)}=\left(\begin{array} {cccccc} 0 & 0 & 0 & 1 & 0 \\ 0 & 1 & 0 & 0 & 0 \\ 0 & 0 & 0 & 1 & 1 \\ 1 & 0 & 0 & 1 & 0 \\ 0 & 1 & 0 & 0 & 1 \end{array}\right)"



"W^{(3)}=\left(\begin{array} {cccccc} 0 & 0 & 0 & 1 & 0 \\ 0 & 1 & 0 & 0 & 0 \\ 0 & 0 & 0 & 1 & 1 \\ 1 & 0 & 0 & 1 & 0 \\ 0 & 1 & 0 & 0 & 1 \end{array}\right)"


"W^{(4)}=\left(\begin{array} {cccccc} 1 & 0 & 0 & 1 & 0 \\ 0 & 1 & 0 & 0 & 0 \\ 1 & 0 & 0 & 1 & 1 \\ 1 & 0 & 0 & 1 & 0 \\ 0 & 1 & 0 & 0 & 1 \end{array}\right)"


Thus, the matrix of transitive closure of "R" is the following:


"M_{R^*}=W^{(5)}=\left(\begin{array} {cccccc} 1 & 0 & 0 & 1 & 0 \\ 0 & 1 & 0 & 0 & 0 \\ 1 & 1 & 0 & 1 & 1 \\ 1 & 0 & 0 & 1 & 0 \\ 0 & 1 & 0 & 0 & 1 \end{array}\right)"



Therefore, the transitive closure of "R" is the following:


"R^* = \{(1,1),(1,4), (2,2),(3,1),(3,2), (3,4), (3,5), (4,1),(4,4), (5,2), (5,5)\}"